# Adaptive Draft-Verification for Efficient Large Language Model Decoding

### v1

@@ Abstract
Large language model (LLM) decoding involves generating a sequence of tokens based on a given context, where each token is predicted one at a time using the model’s learned probabilities. The typical autoregressive decoding method requires a separate forward pass through the model for each token generated, which is computationally inefficient and poses challenges for deploying LLMs in latency-sensitive scenarios. The main limitations of current decoding methods stem from their inefficiencies and resource demands. Existing approaches either necessitate fine-tuning smaller models, which is resource-intensive, or rely on fixed retrieval schemes to construct drafts for the next tokens, which lack adaptability and fail to generalize across different models and contexts. To address these issues, we introduce a novel methodology called ADED  [^1] Project repo: https://anonymous.4open.science/r/ADED-C7D5, which accelerates LLM decoding without requiring fine-tuning. Our approach involves an adaptive draft-verification process that evolves over time to improve efficiency. We utilize a tri-gram matrix-based LLM representation to dynamically approximate the output distribution of the LLM, allowing the model to adjust to changing token probabilities during the decoding process. Additionally, we implement a draft construction mechanism that effectively balances exploration and exploitation, ensuring that the drafts generated are both diverse and close to the true output distribution of the LLM. The importance of this design lies in its ability to optimize the draft distribution adaptively, leading to faster and more accurate decoding. Through extensive experiments on various benchmark datasets and LLM architectures, we demonstrate that ADED significantly accelerates the decoding process while maintaining high accuracy, making it suitable for deployment in a wide range of practical applications.
1 Introduction
@@
Large language model (LLM) decoding involves generating a sequence of tokens based on a given context, where each token is predicted one at a time using the model’s learned probabilities (Brown et al., 2020; Zhang et al., 2022; Touvron et al., 2023a, b). The core mechanism is autoregressive, where each new token is generated conditioned on the previously generated tokens and the given context. This process is crucial for applications like text generation (Li et al., 2024a; Peng et al., 2023; Chang et al., 2023), machine translation (Zhang et al., 2023; Moslem et al., 2023; Hendy et al., 2023), and conversational AI (Shanahan, 2024; Wu et al., 2023; Saka et al., 2023). However, each decoding step involves a forward pass through the model, making the process inherently sequential and computationally expensive. The inefficiencies arise due to the need to reload the model for each token prediction, leading to high computational costs and memory bandwidth usage. This serial nature of decoding is a significant bottleneck, especially for real-time applications (Liu et al., 2023a; Mandvikar, 2023; Antoniol et al., 1994) where latency is critical. Thus, optimizing the decoding speed of LLMs is essential for practical deployment in various real-world scenarios.
@@
Figure 1: Comparison of different LLM decoding strategies: Speculative Decoding Leviathan et al. (2023), Lookahead Fu et al. (2024a), REST He et al. (2023), and our ADED. In Speculative Decoding, a small LLM generates predictions (red blocks) from inputs (blue blocks). Yellow blocks indicating intermediate results obtained from language model. Lookahead uses a large LLM for forward-looking predictions. REST employs a corpus trie for rapid token lookups. ADED integrates Monte Carlo Tree Search with tri-gram (Martin et al., 1998; Golding and Schabes, 1996; Zhu and Rosenfeld, 2001) statistics and recent token history to simulate potential outputs, refining its recommendations over time. ADED’s adaptive approach offers significant advantages in terms of speed and accuracy by continuously evolving its draft constructions, providing more efficient and accurate LLM decoding compared to the fixed or resource-intensive methods used by the others.
@@
Recent research has explored various strategies to mitigate the inefficiencies of LLM decoding. Speculative Decoding (Leviathan et al., 2023; Spector and Re, 2023; Chen et al., 2023) introduces an approach where a smaller, more efficient model generates several token predictions in parallel, which are then verified by the larger target model. This method leverages the efficiency of smaller models to reduce the number of serial forward passes required, achieving substantial speedups without altering the output distribution. Lookahead Decoding Fu et al. (2024a) uses the full context to predict multiple future tokens, creating a buffer that reduces the dependency on sequential processing. REST He et al. (2023) employs a retrieval-based approach where relevant tokens are fetched from a pre-constructed datastore using the current context, forming drafts that are verified by the LLM. These methods can be summarized within the draft-verification pipeline, as shown in Figure 4. Speculative Decoding and Lookahead Decoding both generate draft tokens through predictive models, while REST constructs drafts from retrieved tokens based on the context. In each case, the drafts are then verified by the main LLM, ensuring that the final output adheres to the model’s learned probabilities. Despite their advancements, these approaches face notable limitations. They often require additional training or fine-tuning, which can be resource-intensive. Fixed retrieval schemes lack adaptability, making it challenging to adjust the draft distribution in real-time based on the evolving LLM output. Additionally, these methods may not generalize well across different models and contexts, limiting their effectiveness in dynamic environments.
@@
In this work, our focus is on fine-tuning-free draft-verification to address these limitations. The draft-verification pipeline can be viewed as a rejection sampling procedure where the similarity between the proposal distribution (draft) and the target distribution (LLM output) is crucial for the acceptance rate and convergence speed. Higher similarity results in a higher acceptance rate and faster decoding speed. Very few fine-tuning-free approaches, e.g., REST He et al. (2023), typically use fixed retrieval-based schemes to construct drafts. These schemes lack the adaptability to adjust the draft distribution based on the evolving LLM output distribution, resulting in a persistent gap between the draft and the actual LLM output. This gap reduces the draft acceptance rate and limits the potential for improving decoding speed. To address this issue, we raise the following question:
@@
Question: How to design an adaptive draft construction process that can evolve itself and accurately approximate LLM outputs during decoding?
@@
To introduce adaptability and find drafts that are increasingly close to the LLM output distribution during decoding, we not only need to have an adaptive draft construction pipeline but also need to maintain a balance between exploration and exploitation. This balance ensures that speedups can be achieved by leveraging existing knowledge of draft construction while continuously exploring better draft construction capabilities. To achieve this, we propose a novel methodology called ADED (Adaptive Draft-Verification for Efficient LLM Decoding). ADED incorporates a tri-gram-matrix-based adaptive LLM representative to control the conditional probability distribution of the next token, which can be updated during the decoding process to adjust the draft construction accordingly. To balance exploration and exploitation, we design a draft maker inspired by Monte Carlo Tree Search (MCTS) (Coulom, 2007; Browne et al., 2012; James et al., 2017; Świechowski et al., 2023). This draft maker uses a token preference score to maintain the balance during the search process. The score consists of two parts: the first part is based on the approximate conditional probability distribution of the next token obtained from the LLM representative, reflecting the draft maker’s current knowledge of the LLM output; the second part encourages the draft maker to explore unexplored or less-explored draft spaces. Theoretically, our method can be viewed as a constrained optimization problem to encourage the draft distribution to converge to the LLM output distribution (see Appendix A). Using the token preference score, the draft maker can effectively search the draft space and generate candidate tokens. After the draft construction and verification are completed, the information is fed back to the LLM representative to update its approximation of the LLM output. This feedback loop enriches the draft maker’s knowledge in subsequent rounds of draft-verification, enabling adaptability and self-evolution in the draft construction process.
@@
  • We design a tri-gram matrix-based LLM representation that dynamically approximates the LLM output distribution, enhancing adaptability without the need for fine-tuning. This approach addresses the limitation of fixed retrieval schemes by continuously evolving with the model’s predictions.
@@
  • We develop a draft maker inspired by MCTS, which effectively balances exploration and exploitation to generate high-quality drafts. This mechanism improves decoding speed and accuracy by ensuring that the drafts are closely aligned with the LLM’s output distribution. Our experiments show a 2.5X improvement in decoding speed compared to baselines.
@@
  • Through extensive experiments on various benchmark datasets and LLM architectures, we demonstrate that ADED significantly accelerates the decoding process while maintaining high accuracy. Specifically, we achieve up to a 2.5X speedup in latency and an average acceptance rate improvement of 20% over existing methods.
@@
  • Our method reduces computational overhead and memory usage, making it suitable for deployment in a wide range of practical, real-time applications. Our method’s ability to adapt to evolving LLM outputs and continuously refine draft construction sets it apart from existing, addressing the need for more flexible and dynamic LLM decoding solutions.
2 Methodology
@@
We propose a new fast fine-tuning-free draft-verification LLM decoding method by introducing adaptability into the decoding and learning from LLM. Existing accelerated decoding algorithms either require additional fine-tuning or lack adaptability to LLM’s output distributions, resulting in significant additional cost or insufficient acceleration. To address these issues, we design an adaptive LLM representation based on a tri-gram matrix to adaptively approximate the output distribution of the LLM; develop an MCTS-based draft maker that balances exploration and exploitation for self-evolution towards high-quality drafts; and verify the drafts using tree attention.
2.1 Preliminary: Speculative Decoding & Monte Carlo Tree Search
@@
Speculative decoding is a method to accelerate language model inference by using a smaller auxiliary model to generate a draft sequence, reducing the computational load on the larger model Leviathan et al. (2023). Retrieval-based speculative decoding extends this by incorporating a retrieval system instead of the smaller model, leveraging pre-stored corpus segments for relevant text generation. Monte Carlo Tree Search (MCTS) (Coulom, 2007; Browne et al., 2012; James et al., 2017; Świechowski et al., 2023) is an AI algorithm that optimizes decision-making by balancing exploration and exploitation of future states. It selects nodes for further exploration using a combination of node visit counts and estimated values, aiming to maximize overall outcomes. For a comprehensive discussion of these methods, please refer to Appendix C.
2.2 Adaptive LLM Representative
@@
In order to approximate the output token distribution of the LLM without fine-tuning the small model, we distill linguistic knowledge from a small corpus and construct a tri-gram matrix as an initial representation of the LLM, which allows us to leverage the statistical regularities of language at a granular level. Specifically, we summarize and count the three tokens that appear in the corpus and compute the probability of the third token appearing conditional on the first two tokens. The formula is as defined in Eq. (1):
$P(w_{i}|w_{i-2},w_{i-1})=\frac{C(w_{i-2},w_{i-1},w_{i})}{C(w_{i-2},w_{i-1})},\vspace{-2mm}$ (1)
@@
where $P(w_{i}|w_{i-2},w_{i-1})$ is the conditional probability of a word $w_{i}$ given the two preceding words $w_{i-2}$ and $w_{i-1}$, $C(w_{i-2},w_{i-1},w_{i})$ is the count of the tri-gram occurrence in the corpus, and $C(w_{i-2},w_{i-1})$ is the count of the preceding bi-gram Mori et al. (1998).
@@
In this way, we can obtain a good initial LLM representative at a much lower cost, which can generate an approximate distribution of the next token based on the previous tokens. This LLM representative will collaborate with our draft maker to generate drafts and get feedback to update the tri-gram matrix for adaptability and self-evolution. Please see Section 2.3 for more details.
2.3 Draft Maker and Self-Evolution
@@
With the help of the LLM representative, we further propose a draft maker that balances exploration and exploitation while searching for candidate drafts that are closer to the LLM output. On the one hand, our draft maker leverages the conditional probabilities from the LLM representative, which include current knowledge of the LLM output. On the other hand, our draft maker is encouraged to search more in the unexplored or less explored draft space to find better draft candidates. Then, with feedback from the LLM output, the LLM representative can update its understanding of the LLM output, improve the draft maker’s search, and achieve self-evolution. Details are provided below.
@@
Draft Search Score: Given the initial tokens, we exploit Monte Carlo Tree Search (MCTS) Coulom (2007) to guide the search process of the drafts of the next tokens, where we prioritize candidate tokens according to the conditional probability from the tri-gram matrix-based LLM representative and the node visitation counts during the tree search. Our scores play a key role in balancing exploration and utilization during Monte Carlo tree search and is defined as Eq. (2). This is a kind of PUTC Score (Rosin, 2011; Silver et al., 2017). More specifically, $Q(s,a)$ assesses the quality of taking action $a$ in state $s$, while $P(s,a)$ represents the prior probability of selecting action $a$ in state $s$. The term $N(s,a)$ denotes the number of times the action $a$ has been taken from state $s$, and $N(s,b)$ sums the counts for all actions from state $s$. The constants $c_{1}$ and $c_{2}$ adjust the balance between exploration and exploitation, improving the decision-making process in draft construction. This formula ensures that our draft choices are contextually appropriate and optimizes the robustness and coherence of text generation.
$\max{Q(s,a)+P(s,a)\cdot\frac{\sqrt{\sum_{b}N(s,b)}}{1+N(s,a)}}(c_{1}+\log(\frac{\sum_{b}N(s,b)+c_{2}+1}{c_{2}})).$ (2)
Figure 2: Self-Evolution strategy transfer. This diagram shows how the LLM updates its tri-gram probabilities based on recent outputs, feeding these probabilities into the MCTS. This dynamic update enhances the generation of coherent sequences by evolving the search strategy over time.
@@
Self-Evolution Strategy Transfer: Based on the final search score obtained during the Monte Carlo tree search, we can construct draft candidates and verify them to get the final decoding output (please see Section 2.4) and feed it back for self-evolution. This final output decoding represents LLM’s output distribution, which would be a good learning material for the LLM representative. Therefore, we feed this knowledge into the LLM representative in order to obtain updated conditional probability distributions, thus providing the draft maker with more accurate and exploitable knowledge, which is illustrated in Figure 2. Specifically, this technique operates by first extracting tri-grams from recent outputs of the LLM. Each tri-gram’s frequency is then used to update its probability as potential outputs. These adjusted probabilities are fed into the MCTS as part of the policy network, influencing the selection phase of the tree search. In the context of MCTS, the updated tri-gram probabilities essentially serve as a dynamic policy guide, enhancing the model’s ability to generate contextually relevant and coherent sequences. By incorporating learned tri-gram probabilities into the tree search algorithm, we effectively create a feedback loop where the search strategy itself evolves over time. This strategy adjustment is executed by recalibrating the exploration-exploitation balance based on the empirical data derived from the model’s own outputs.
2.4 Draft Construction and Verification
@@
To validate the draft sequences, it is noted that many have common starting segments that can cause redundant recalculations in the Transformer layers if not managed correctly. To address the issue, a pseudo-sequence that guarantees that each draft is a sub-sequence and that any common prefix appears only once is created He et al. (2023). We also use a specific attention mask for each attention layer, called tree attention (Miao et al., 2023; Cai et al., 2024). This mask aligns the computations for each token with its dependencies according to the original draft sequence, preserving the draft’s contextual integrity and preventing unnecessary computations. The approval of drafts relies on a comparison with the conditional distribution from the LLM. At each position, new tokens are sampled and compared to the draft tokens. If a sampled token corresponds to the draft token, it is approved; otherwise, the draft is discarded from that point. This selective approval ensures that the output sequence aligns with what would be produced by a typical autoregressive process, thus upholding the authenticity of the generated text.
3 Theoretical Insight: Why ADED uses MCTS
@@
In this section, we explore the theoretical parallels between the Monte Carlo Tree Search (MCTS) algorithm used in our ADED framework and the inference mechanisms of large language models (LLMs) to demonstrate the use of MCTS and the self-evolution of ADED. We show that draft search in ADED using MCTS can be viewed as a form of policy optimization, and that the inference mechanism of LLM can be viewed as a similar form of penalty optimization.
@@
where $n(x,a)$ represents the visit count for action $a$ in state $x$. Then, the action selection in MCTS can be written as selecting the action $a^{*}$:
@@
Following (Grill et al., 2020), we use $q\in\mathcal{R}^{|A|}$ to denote the vector of Q-function $Q(x,a)$. With proper choice of hyper-parameters, the MCTS algorithm can be viewed as searching for the optimum solution to a policy optimization problem (Grill et al., 2020) as below:
@@
LLM Inference Mechanism: Large language models, particularly those based on the Transformer architecture, generate text by predicting the probability distribution of the next token given the previous tokens. During inference, the model maximizes the log-likelihood of the observed data, which is equivalent to minimizing the cross-entropy loss:
$\mathcal{L}(\theta)=-\sum_{t=1}^{T}\log P(w_{t}\mid w_{1:t-1};\theta),$ (6)

### v2

@@ Abstract
Large language model (LLM) decoding involves generating a sequence of tokens based on a given context, where each token is predicted one at a time using the model’s learned probabilities. The typical autoregressive decoding method requires a separate forward pass through the model for each token generated, which is computationally inefficient and poses challenges for deploying LLMs in latency-sensitive scenarios. The main limitations of current decoding methods stem from their inefficiencies and resource demands. Existing approaches either necessitate fine-tuning smaller models, which is resource-intensive, or relying on fixed retrieval schemes to construct drafts for the next tokens, which lack adaptability and fail to generalize across different models and contexts. To address these issues, we introduce a novel methodology called ADED  [^1] Project repo: https://anonymous.4open.science/r/ADED-C7D5, which accelerates LLM decoding without requiring fine-tuning. Our approach involves an adaptive draft-verification process that evolves over time to improve efficiency. We utilize a tri-gram matrix-based LLM representation to dynamically approximate the output distribution of the LLM, allowing the model to adjust to changing token probabilities during the decoding process. Additionally, we implement a draft construction mechanism that effectively balances exploration and exploitation, ensuring that the drafts generated are both diverse and close to the true output distribution of the LLM. The importance of this design lies in its ability to optimize the draft distribution adaptively, leading to faster and more accurate decoding. Through extensive experiments on various benchmark datasets and LLM architectures, we demonstrate that ADED accelerates the decoding process while maintaining high accuracy, making it suitable for deployment in a wide range of practical applications.
Introduction
@@
Large language model (LLM) decoding involves generating a sequence of tokens based on a given context, where each token is predicted one at a time using the model’s learned probabilities (Brown et al. 2020; Zhang et al. 2022; Touvron et al. 2023a, b). The core mechanism is autoregressive, where each new token is generated conditioned on the previously generated tokens and the given context. This process is crucial for applications like text generation (Li et al. 2024a; Peng et al. 2023; Chang et al. 2023), machine translation (Zhang, Haddow, and Birch 2023; Moslem et al. 2023; Hendy et al. 2023), and conversational AI (Shanahan 2024; Wu et al. 2023; Saka et al. 2023). However, each decoding step involves a forward pass through the model, making the process inherently sequential and computationally expensive. The inefficiencies arise due to the need to reload the model for each token prediction, leading to high computational costs and memory bandwidth usage. This serial nature of decoding is a significant bottleneck, especially for real-time applications (Liu et al. 2023a; Mandvikar 2023; Antoniol et al. 1994) where latency is critical. Thus, optimizing the decoding speed of LLMs is essential for practical deployment.
@@
Figure 1: Comparison of different LLM decoding strategies. In Speculative Decoding, a small LLM generates predictions (red blocks) from inputs (blue blocks). Yellow blocks indicating intermediate results obtained from language model. Lookahead uses a large LLM for forward-looking predictions. REST employs a corpus trie for rapid token lookups. ADED integrates Monte Carlo Tree Search with tri-gram statistics and recent token history to simulate potential outputs, refining its recommendations over time. ADED’s adaptive approach offers advantages in terms of speed and accuracy by continuously evolving its draft constructions.
@@
Recent research has explored various strategies to mitigate the inefficiencies of LLM decoding. Speculative Decoding (Leviathan, Kalman, and Matias 2023; Spector and Re 2023; Chen et al. 2023) introduces an approach where a smaller, more efficient model generates several token predictions in parallel, which are then verified by the larger target model. This method leverages the efficiency of smaller models to reduce the number of serial forward passes required, achieving substantial speedups without altering the output distribution. Lookahead Decoding (Fu et al. 2024a) uses the full context to predict multiple future tokens, creating a buffer that reduces the dependency on sequential processing. REST (He et al. 2024) employs a retrieval-based approach where relevant tokens are fetched from a pre-constructed datastore using the current context, forming drafts that are verified by the LLM. These methods can be summarized within the draft-verification pipeline, as shown in Figure 1. Speculative Decoding and Lookahead Decoding both generate draft tokens through predictive models, while REST constructs drafts from retrieved tokens based on the context. In each case, the drafts are then verified by the main LLM, ensuring that the final output adheres to the model’s learned probabilities. Despite their advancements, these approaches face notable limitations. They often require additional training or fine-tuning, which can be resource-intensive. Fixed retrieval schemes lack adaptability, making it challenging to adjust the draft distribution in real-time based on the evolving LLM output. Additionally, these methods may not generalize well across different models and contexts, limiting their effectiveness in dynamic environments.
@@
In this work, our focus is on fine-tuning-free draft-verification to address these limitations. The draft-verification pipeline can be viewed as a rejection sampling procedure where the similarity between the proposal distribution (draft) and the target distribution (LLM output) is crucial for the acceptance rate and convergence speed. Higher similarity results in a higher acceptance rate and faster decoding speed. Very few fine-tuning-free approaches, e.g., REST (He et al. 2024), typically use fixed retrieval-based schemes to construct drafts. These schemes lack the adaptability to adjust the draft distribution based on the evolving LLM output distribution, resulting in a persistent gap between the draft and the actual LLM output. This gap reduces the draft acceptance rate and limits the potential for improving decoding speed. To address this issue, we raise the following question:
@@
Research Question: How to design an adaptive draft construction process that can evolve itself and accurately approximate LLM outputs during decoding?
@@
To introduce adaptability and find drafts that are increasingly close to the LLM output distribution during decoding, we not only need to have an adaptive draft construction pipeline but also need to maintain a balance between exploration and exploitation. This balance ensures that speedups can be achieved by leveraging existing knowledge of draft construction while continuously exploring better draft construction capabilities. To achieve this, we propose a novel methodology called ADED (Adaptive Draft-Verification for Efficient LLM Decoding). ADED incorporates a tri-gram-matrix-based adaptive LLM representative to control the conditional probability distribution of the next token, which can be updated during the decoding process to adjust the draft construction accordingly. To balance exploration and exploitation, we design a draft maker inspired by Monte Carlo Tree Search (MCTS) (Coulom 2007; Browne et al. 2012; James, Konidaris, and Rosman 2017; Świechowski et al. 2023). This draft maker uses a token preference score to maintain the balance during the search process. The score consists of two parts: the first part is based on the approximate conditional probability distribution of the next token obtained from the LLM representative, reflecting the draft maker’s current knowledge of the LLM output; the second part encourages the draft maker to explore unexplored or less-explored draft spaces. Theoretically, we show that our method can be viewed as a constrained optimization problem to encourage the draft distribution to converge to the LLM output distribution. Using the token preference score, the draft maker can effectively search the draft space and generate candidate tokens. After the draft construction and verification are completed, the information is fed back to the LLM representative to update its approximation of the LLM output. This feedback loop enriches the draft maker’s knowledge in subsequent rounds of draft-verification, enabling adaptability and self-improvement in the draft construction process.
@@
  • We design a tri-gram matrix-based representation that dynamically approximates the LLM output distribution, enhancing adaptability without the need for fine-tuning. It addresses the limitation of fixed retrieval schemes by continuously evolving with the model’s predictions.
@@
  • We develop a draft maker that effectively balances exploration and exploitation to generate high-quality drafts. This mechanism improves decoding speed and accuracy by ensuring that the drafts are closely aligned with the LLM’s output distribution. Our experiments show a 2.5X improvement in decoding speed compared to baselines.
@@
  • Through extensive experiments on various benchmark datasets and LLM architectures, we demonstrate that ADED successfully accelerates the decoding process while maintaining high accuracy. Specifically, we achieve up to a 2.5X speedup in latency and an average acceptance rate improvement of 20% over existing methods.
@@
  • Our method’s ability to adapt to evolving LLM outputs and continuously refine draft construction sets it apart from existing ones, addressing the need for more flexible and dynamic decoding solutions.
Methodology
Figure 2: This figure illustrates the data processing workflow of ADED. Initially, the input tokens undergo preprocessing to calculate their tri-grams, which serve to update the tri-gram matrix. Subsequently, the updated matrix, in conjunction with the last two tokens of the input, is used to retrieve potential token sequences. These sequences are ranked, and the top-k sequences are selected, and then appended to the original input. Finally, these extended sequences are inputted into the Large Language Model for prediction.
@@
We propose a new fast fine-tuning-free draft-verification LLM decoding method by introducing adaptability into the decoding and learning from LLM, which is illustrated in Figure 2. Existing accelerated decoding algorithms either require additional fine-tuning or lack adaptability to LLM’s output distributions, resulting in additional cost or insufficient acceleration. To address these issues, we design an adaptive LLM representation based on a tri-gram matrix to adaptively approximate the output distribution of the LLM, develop a draft maker that balances exploration and exploitation for self-improvement towards high-quality drafts, and verify the drafts using tree attention.
Preliminary
@@
Speculative decoding is a method to accelerate language model inference by using a smaller auxiliary model to generate a draft sequence, reducing the computational load on the larger model (Leviathan, Kalman, and Matias 2023). Retrieval-based speculative decoding extends this by incorporating a retrieval system instead of the smaller model, leveraging pre-stored corpus segments for relevant text generation. Monte Carlo Tree Search (MCTS(Coulom 2007; Browne et al. 2012; James, Konidaris, and Rosman 2017; Świechowski et al. 2023) is an algorithm that optimizes decision-making by balancing exploration and exploitation of future states. It selects nodes for further exploration using a combination of node visit counts and estimated values, aiming to maximize overall outcomes. For a comprehensive discussion of these methods, please refer to Appendix E.
Adaptive LLM Representative
@@
To approximate the output token distribution of the LLM without fine-tuning the small model, we distill linguistic knowledge from a small corpus and construct a tri-gram matrix as an initial representation of the LLM, which allows us to leverage the statistical regularities of language at a granular level. Specifically, we summarize and count each set of three tokens that appear in the corpus and compute the probability of the third token appearing conditional on the first two tokens. The formula is defined in Eq. (1):
$P(w_{i}|w_{i-2},w_{i-1})=\frac{C(w_{i-2},w_{i-1},w_{i})}{C(w_{i-2},w_{i-1})},$ (1)
@@
where $P(w_{i}|w_{i-2},w_{i-1})$ is the conditional probability of a word $w_{i}$ given the two preceding words $w_{i-2}$ and $w_{i-1}$, $C(w_{i-2},w_{i-1},w_{i})$ is the count of the tri-gram occurrence in the corpus, and $C(w_{i-2},w_{i-1})$ is the count of the preceding bi-gram (Mori, Nishimura, and Itoh 1998).
@@
In this way, we can obtain a good initial LLM representative at a much lower cost, which can generate an approximate distribution of the next token based on the previous tokens. This LLM representative will collaborate with our draft maker to generate drafts and get feedback to update the tri-gram matrix for adaptability and self-improvement. Please see Section 2.3 Draft Maker and Self-Improvement for more details.
Draft Maker and Self-Improvement
@@
With the help of the LLM representative, we further propose a draft maker that balances exploration and exploitation while searching for candidate drafts that are closer to the LLM output. On the one hand, the draft maker leverages the conditional probabilities from the LLM representative, which include current knowledge of the LLM output. On the other hand, the draft maker is encouraged to search more in the unexplored or less explored draft space to find better draft candidates. Then, with the feedback from the LLM output, the LLM representative can update its understanding of the LLM output, improve the draft maker’s search, and achieve self-improvement. Details are provided below.
@@
Draft Search Score: Given the initial tokens, we exploit Monte Carlo Tree Search (MCTS(Coulom 2007) to guide the search process of the drafts of the next tokens, where we prioritize candidate tokens according to the conditional probability from the tri-gram matrix-based LLM representative and the node visitation counts during the tree search. The score plays a key role in balancing exploration and utilization during the Monte Carlo tree search and is defined as Eq. (2).
$\text{PUCT}(s,a)=Q(s,a)+E\cdot P(s,a)\cdot\frac{\sqrt{\sum_{b}N(s,b)}}{1+N(s,a)}.$ (2)
The score design is motivated by PUCT Score (Rosin 2011; Silver et al. 2017). In particular, $Q(s,a)$ assesses the quality of taking action $a$ in state $s$, while $P(s,a)$ represents the prior probability of selecting action $a$ in state $s$. The term $N(s,a)$ denotes the number of times the action $a$ has been taken from state $s$, and $\sum_{b}N(s,b)$ sums the counts for all actions from state $s$. Eq. (3) plays a critical role in determining the balance between exploration and exploitation within the MCTS framework.
$E=C_{1}+\log\left(\frac{\sum_{b}N(s,b)+C_{2}+1}{C_{2}}\right),$ (3)
The constant $C_{1}$ acts as a base level adjustment, while $C_{2}$ modulates the logarithmic term to scale the exploration factor dynamically based on the total visitation counts. This formula ensures that our draft choices are contextually appropriate and optimizes the robustness and coherence of text generation.
@@
Self-Improvement Strategy Transfer: Based on the final search score obtained during the search, we can construct draft candidates and verify them to get the final decoding output (please see Section 2.4 Draft Construction and Verification) and feed it back for self-improvement. This final output decoding represents LLM’s output distribution, which would be a good learning material for the LLM representative. Therefore, we feed this knowledge into the LLM representative in order to obtain updated conditional probability distributions, thus providing the draft maker with more accurate and exploitable knowledge, which is illustrated in Figure 2. Specifically, this technique operates by first extracting tri-grams from recent outputs of the LLM. Each tri-gram’s frequency is then used to update its probability as potential outputs. These adjusted probabilities are fed into the MCTS as part of the policy network, influencing the selection phase of the tree search. The updated tri-gram probabilities essentially serve as a dynamic policy guide, enhancing the model’s ability to generate contextually relevant and coherent sequences. By incorporating learned tri-gram probabilities into the tree search algorithm, we effectively create a feedback loop where the search strategy itself evolves over time. This strategy adjustment is executed by recalibrating the exploration-exploitation balance based on the empirical data derived from the model’s own outputs.
Draft Construction and Verification
@@
It is important to note that candidate drafts generated by the draft maker often have common starting segments that can cause redundant recalculations in the Transformer layers if not managed correctly. To address the issue, a pseudo-sequence that guarantees that each draft is a sub-sequence and that any common prefix appears only once is created (He et al. 2024). Motivated by this observation, we use a specific attention mask for each attention layer, called tree attention (Miao et al. 2023; Cai et al. 2024). This mask aligns the computations for each token with its dependencies according to the original draft sequence, preserving the draft’s contextual integrity and preventing unnecessary computations. The approval of drafts relies on a comparison with the conditional distribution from the LLM. At each position, new tokens are sampled and compared to the draft tokens. If a sampled token corresponds to the draft token, it is approved; otherwise, the draft is discarded from that point. This selective approval ensures that the output sequence aligns with what would be produced by a typical autoregressive process, thus upholding the authenticity of the generated text.
Theoretical Insight: Why ADED uses MCTS
@@
In this section, we provide a theoretical justification for the design of the ADED method. We show that the draft search in ADED using MCTS can be viewed as a form of policy optimization, while the inference mechanism of LLMs can be viewed as a similar form of penalty optimization.
@@
where $n(x,a)$ represents the visit count for action $a$ in state $x$, and $|A|$ represents the total number of possible actions at state $x$. Then, the action selection in MCTS can be written as selecting the action $a^{*}$:
@@
Following (Grill et al. 2020), we use $q\in\mathcal{R}^{|A|}$ to denote the vector of Q-function $Q(x,a)$. With proper choice of hyper-parameters, the MCTS algorithm can be viewed as searching for the optimum solution to a policy optimization problem (Grill et al. 2020) as below:
@@
LLM Inference Mechanism: Large language models, particularly those based on the Transformer architecture, generate text by predicting the probability distribution of the next token given the previous tokens. During training, the model maximizes the log-likelihood of the observed data, which is equivalent to minimizing the cross-entropy loss:
$\mathcal{L}(\theta)=-\sum_{t=1}^{T}\log P(w_{t}\mid w_{1:t-1};\theta)+\frac{\lambda}{2}\|\theta\|_{2}^{2},$ (7)